OpenStudy (anonymous):

Each of two urns contain green balls and red balls. Urn 1 contains 8 green balls and 12 red balls. Urn 2 contains 5 green balls and 8 red balls. If a ball is drawn from each urn, what is P(red and red)? a. 79/65 b. 24/65 c. 20/33 d. 2/13

OpenStudy (anonymous):

ok u have to find seperately two probibailities, one for Urn1 red ball and 2nd for Urn2 red ball.. then u you have ti take there AND and that will be ur answer

OpenStudy (anonymous):

did u get it ?

OpenStudy (anonymous):

idk how?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well first we find the Probibality for Urn1 red ball Urn contains 8 green and 12 red total balls are 12+8=20 the so when we pick one ball from it the chance of being red is its probibality it will be 12/20 got it till now its for Urn1

OpenStudy (anonymous):

P(red for urn1)= 12/20 = 3/5 P(red for urn2)= 8/13 P(red AND red)= (3/5)x(8/13) = 24/65

OpenStudy (anonymous):

@CO_oLBoY try to explain her .. do not make her dependant plz .. dnt just give answers . she needs explanation

OpenStudy (anonymous):

ok same for Urn2 u have total 5+8=13 balls 8 are red so its probibality is 8/13 right

OpenStudy (anonymous):

i didnt just gave the ans @khurramshahzad

OpenStudy (anonymous):

now u have to take there and /.. u knw how to take and .. just multiply both probibalities .. its the end result .. as @CO_oLBoY shown above

OpenStudy (anonymous):

ok bro @CO_oLBoY

OpenStudy (anonymous):

@khurramshahzad asal me sab ko ans hi chahye hota to thora sa explain kr k ans de dya kro

OpenStudy (anonymous):

well if they just need answer .. y are v here are they paying us or something like that .. we just help to make them learn.. if i told her this way .. it will surely make her do it next time by her self .. as i think man ..